# Each of those eight coalition logics is also determined by six other kinds of models

### v1

@@ 8 Each of those eight coalition logics is determined by these six kinds of models, too
  • Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent. Let $\chi\in\mathtt{ST}^{\prime}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\chi)$ and $\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{D}}(\chi)$. Then, $\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}^{\prime}_{\mathrm{D}}(\chi,\sigma_{\mathrm{D}})\neq\emptyset$.
    Then, there is $s\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$. Then $s\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then $s=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$. Similarly, we can show $\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(\mathsf{last}(\chi))$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi))$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. Let $u^{\prime}\in\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$.
    Then $u^{\prime}\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma^{\prime}_{\mathtt{AG}})$ for some $\sigma^{\prime}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\subseteq\sigma^{\prime}_{\mathtt{AG}}$.
    Then $\chi-\sigma^{\prime}_{\mathtt{AG}}-u^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma^{\prime}_{\mathtt{AG}})$.
    Then, $\chi-\sigma^{\prime}_{\mathtt{AG}}-u^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\chi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})$. Then, $\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\chi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\chi)$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is independent.
@@
  • Assume $(\mathtt{G}\textbf{-}\mathtt{AM},w)$ is deterministic. Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is not deterministic. Then, there is $\chi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(\chi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ is not a singleton. Let $s,s^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ where $s\neq s^{\prime}$. Let $s=\chi-\sigma_{\mathtt{AG}}-u$ and $s^{\prime}=\chi-\sigma_{\mathtt{AG}}-u^{\prime}$ where $u\neq u^{\prime}$. Then, $u,u^{\prime}\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$. Then, there is a contradiction. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is deterministic.
@@
It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is a tree-like model.
We show for all $\chi$, and $\phi$, $\mathtt{G}\textbf{-}\mathtt{AM},w\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w\Vdash\phi$.
We put an induction on $\phi$. We consider only the cases $\phi=p$ and $\phi=\langle[\mathrm{C}]\rangle\psi$, and skip other cases.
  • $\phi$ is a proposition letter $p$.
    $\Leftarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash p$. Then, $p\in\mathsf{label}^{\prime}(\chi)$. Then, $p\in\mathsf{label}(\mathsf{last}(\chi))\}$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$.
    $\Rightarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash p$. Then, $p\in\mathsf{label}(\mathsf{last}(\chi))\}$. Then, $p\in\mathsf{label}^{\prime}(\chi)$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$.
  • $\phi$ is $\langle[\mathrm{C}]\rangle\psi$.
@@
    $\Leftarrow$ Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\chi\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\chi)$ such that for all $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
    We are done if we can show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
    Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})\neq\emptyset$. Let $s\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$. Then $s\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then $s=\chi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$. Then, $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$.
    Let $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})\text{and}$ $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then $y=\chi-\sigma_{\mathtt{AG}}-t\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$.
    Then, $y\in\mathsf{out}^{\prime}(\chi,\sigma_{\mathrm{C}})$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$. By the inductive hyphothesis, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(y)\Vdash\psi$. That is $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
    $\Rightarrow$
@@
    Assume $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\phi$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},\mathsf{last}(\chi)\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\mathsf{last}(\chi))$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
    We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$ and for all $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
    Note $\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})\neq\emptyset$. Let $t\in\mathsf{out}_{\mathrm{C}}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$. Then $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then $y=\chi-\sigma_{\mathtt{AG}}-t\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})$. Then $y\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathrm{C}})$. Then $\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathrm{C}})$ is not empty. Then $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\chi)$.
    Let $y\in\mathsf{out}^{\prime}_{\mathrm{C}}(\chi,\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $y\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\chi,\sigma_{\mathtt{AG}})\text{ and}$ $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$.
    Then, there is $t\in\mathsf{out}_{\mathtt{AG}}(\mathsf{last}(\chi),\sigma_{\mathtt{AG}})$ such that $t=\mathsf{last}(y)$.
    Then, $t\in\mathsf{out}(\mathsf{last}(\chi),\sigma_{\mathrm{C}})$. Then, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},y\Vdash\psi$.
@@
For every $\mathtt{X}\in\mathrm{ES}$ and $\phi\in\Phi$, the following statements are equivalent:
@@
By Theorem 8 and Theorem 1, (2), and (3) are equivalent.
By Theorem 14 and Theorem 1, (4), and (5) are equivalent.
By Theorem 20 and Theorem 1, (6), and (7) are equivalent.
It suffices to show that (1), (2), (4) and (6) are equivalent. By Theorem 4, (1) implies (2). By Theorem 10, (2) implies (4). By Theorem 16, (4) implies (6). Suppose (6). Assume (1) does not hold. Then $\phi$ is false at a pointed grand-coalition-first action model $(\mathtt{G}\textbf{-}\mathtt{AM},w)$. Then, by Theorem 21, there is a tree-like grand-coalition-first model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ such that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)\nVdash\phi$. We have a contradiction. Then, (1).
9 Concluding remarks
Grand-coalition-first action models are those action models in which the outcome function for the grand coalition determines the outcome functions for other coalitions. Single-coalition-first action models are those action models in which the outcome functions for single coalitions determine the outcome functions for other coalitions. This class of models can be $z$-represented by the class of single-coalition-first neighborhood models. Clear grand-coalition-first action models are those single-coalition-first action models where different joint actions of a coalition have different outcomes. This class of models can be $z$-represented by the class of clear single-coalition-first neighborhood models. Tree-like single-coalition-first action models are those single-coalition-first action models where every state has a determined history. This class of models can be $z$-represented by the class of tree-like single-coalition-first neighborhood models. These kinds of models have different motivations.
Every kind of model has eight subclasses, determined by which of the properties seriality, independence of agents, and determinism hold. In this paper, we show that the subclasses of models that meet the same properties determine the same coalition logic.

### v2

@@ 7 Each of those eight coalition logics is determined by these six kinds of models, too
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is independent. Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$, and $\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{D}}(\pi)$. To show $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is independent, it suffices to show $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\pi)$. Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}^{\prime}_{\mathrm{D}}(\pi,\sigma_{\mathrm{D}})\neq\emptyset$. By the claim (1), $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$ and $\mathsf{out}_{\mathrm{D}}(\pi^{l},\sigma_{\mathrm{D}})\neq\emptyset$. Then $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ and $\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{D}}(\pi^{l})$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}_{\mathrm{C}\cup\mathrm{D}}(\pi^{l})$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(\pi^{l},\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. By the claim (2), $\mathsf{out}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\pi,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}\in\mathtt{av}^{\prime}_{\mathrm{C}\cup\mathrm{D}}(\pi)$.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic. Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is not deterministic. Then, there is $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(\pi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ is not a singleton. Let $\pi^{\prime},\pi^{\prime\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ such that $\pi^{\prime}\neq\pi^{\prime\prime}$. Let $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime}$ and $\pi^{\prime\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime\prime}$. Note $t^{\prime}\neq t^{\prime\prime}$. Then, $t^{\prime},t^{\prime\prime}\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$. We have a contradiction.
@@
It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is tree-like.
We show for all $\pi\in\mathtt{ST}^{\prime}$ and $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\phi$, which implies for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$. We put an induction on $\phi$. We consider only the case $\phi=\langle[\mathrm{C}]\rangle\psi$ and skip others.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ such that for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi)$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$. By the claim (1), $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$. Let $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. By the claim (2), there is a $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime l}=t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Note $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ and for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. Note $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$. By the claim (2), $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$. Then $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$. Let $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. By the claim (1), $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Note $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{\prime l}\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
@@
For all $\mathtt{X}\in\mathrm{ES}$ and $\phi\in\Phi$, the following statements are equivalent:
@@
By Theorem 8 and Theorem 1, (1) and (2) are equivalent. By Theorem 14 and Theorem 1, (3) and (4) are equivalent. By Theorem 20 and Theorem 1, (5) and (6) are equivalent. It suffices to show that (0), (1), (3) and (5) are equivalent. By Theorem 4, (0) implies (1). By Theorem 10, (1) implies (3). By Theorem 16, (3) implies (5). By Theorem 21, (5) implies (1).
8 Conclusion
Li and Ju [9, 8] proposed eight coalition logics based on grand-coalition-first action models. In this work, we show that each of them is determined by six kinds of models, that is, single-coalition-first action models, single-coalition-first neighborhood models, clear grand-coalition-first action models, clear single-coalition-first neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first neighborhood models.